# CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs

### v2

@@ 4.1 Experimental Setup
NMTE 0.477 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
$\text{CoCoA}_{NMTE}$ 0.548 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
Table 1: Results for Evaluated Sequence – Best Sample: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Similarity Function.
To measure the similarity, we use the RoBERTa-large cross-encoder model, fine-tuned on the Semantic Textual Similarity benchmark dataset (Liu et al., 2019; Reimers and Gurevych, 2019; Cer et al., 2017). This model is widely regarded as one of the most reliable and commonly used approaches for evaluating sentence similarity. The cross-encoder processes two sequences jointly and directly outputs a similarity score ranging from 0 to 1, providing a nuanced measure.
Appendix B.2 contains comparative experiments with cross-encoder and other choices of the similarity function, substantiating this choice.
Baselines.
We compare the performance of the proposed method against a diverse set of baseline and state-of-the-art UQ scores, including confidence-based, consistency-based, and hybrid approaches. Information-based methods include Maximum Sequence Probability (MSP), Perplexity (PPL), Mean Token Entropy (MTE), Monte Carlo Sequence Entropy (MCSE), and Monte Carlo Normalized Sequence Entropy (MCNSE). Consistency-based methods include the Degree Matrix (DegMat) and the Sum of Eigenvalues of the Graph Laplacian (EigValLaplacian). Finally, Hybrid methods include Semantic Entropy and SAR. All formulations for these baselines can be found in Appendix C.
Evaluation measure.
As our evaluation measure, we choose the Prediction Rejection Ratio (PRR), which measures the effectiveness of the uncertainty scores for identifying high-quality predictions (Malinin and Gales, 2021). PRR operates by progressively rejecting predictions with uncertainty scores above a threshold $a$ and observing how the average quality $Q(f(\mathbf{x}_{i}),\mathbf{y}_{i})$ of the remaining predictions changes. The metric is calculated as the ratio of two areas: the area between the Prediction Rejection (PR) curves for the evaluated uncertainty score and a random baseline, and the area between the oracle (the ideal uncertainty score that perfectly ranks instances by quality) and the random baseline. Formally, PRR is defined as follows:
Higher PRR values indicate better alignment of uncertainty scores with prediction quality, approaching the performance of an oracle. To ensure practical applicability, we compute PRR only up to a rejection threshold of 50%, preventing cases where excessive rejection artificially inflates quality measures. Figure 3 gives a visual representation of the PRR calculation, highlighting the relationship between the uncertainty threshold and the quality measures.
Metric Llama Mistral Falcon
@@
MCSE 0.310 0.323 0.033 0.389 0.304 0.007 0.414 0.317 0.159
MCNSE 0.309 0.393 0.022 0.384 0.410 0.009 0.405 0.422 0.108
SAR 0.414 0.455 0.077 0.462 0.435 0.094 0.481 0.458 0.144
DegMat 0.406 0.302 0.081 0.423 0.305 0.137 0.483 0.353 0.201
MSP 0.409 0.399 0.328 0.475 0.383 0.287 0.475 0.356 0.201
PPL 0.381 0.386 0.369 0.424 0.427 0.204 0.456 0.450 0.155
NMTE 0.353 0.382 0.357 0.417 0.438 0.182 0.456 0.473 0.152
$\text{CoCoA}_{NMTE}$ 0.447 $\uparrow$ 0.478 $\uparrow$ 0.380 $\uparrow$ 0.492 $\uparrow$ 0.469 $\uparrow$ 0.288 $\uparrow$ 0.527 $\uparrow$ 0.508 $\uparrow$ 0.228 $\uparrow$
Quality Measures.
The Predictive Rejection Ratio (PRR) requires an appropriate quality measure for each specific task to effectively evaluate the model output. For question-answering tasks, we use Accuracy to directly evaluate whether the generated answers match the ground truth in short-form QA tasks (e.g., MMLU), and we use the AlignScore between correct answer and generated sequence for assessing the performance for long-form QA tasks (Zha et al., 2023). For summarization tasks, we use AlignScore to measure the alignment between the output summary and the input document. It serves as a quality indicator by evaluating the relevance and the overlap between the generated content with the source text. For translation tasks, we use COMET, as it captures both semantic adequacy and fluency, ensuring that translations are accurate and linguistically appropriate (Rei et al., 2020).
Generation Setup.
We discuss the generation parameters, the decoding strategy and sample selection procedure in depth in Appendix D. In short, we report evaluation results in two distinct setups: greedy decoding and stochastic sampling with focus on the most probable sequence among the generated outputs (best-sample). These two setups offer the highest-quality outputs and are the most reasonable generation approaches in practice.
4.2 Results
Main results.
Tables 1 and 2 show the PRR scores under the best‐sample and greedy generation setup. We calculate a single representative PRR for each task – question answering, neural machine translation (NMT), and summarization (SUM) – by averaging the results across all relevant datasets (e.g., TriviaQA, MMLU, CoQA, GSM8k for QA). This aggregated score provides a concise measure of the performance for each model for each task. Detailed results for each dataset separately can be found in Appendix D.
We can see that our CoCoA methods are the best across all tasks and models in our experiments. They outperform existing consistency‐based and hybrid state‐of‐the‐art approaches, like Semantic Entropy and SAR. In addition, our proposed CoCoA approach consistently surpasses the baseline UE metrics: for example, $\texttt{CoCoA}_{PPL}$ outperforms standard Perplexity, illustrating the advantage of combining token‐level confidence with semantic consistency. This pattern holds for other information‐based metrics as well, demonstrating that using the consistency between multiple sampled outputs reliably enhances uncertainty quantification.
Ablation study.
As a part of our ablation study (see Appendix B.1), we evaluate the performance of the average dissimilarity component ($u_{*}^{\text{cons}}$) independently to assess its effectiveness as a standalone uncertainty measure and to investigate whether it could potentially outweigh the contribution of the information-based component ($u_{*}^{\text{info}}$) in the enriched uncertainty measure. This evaluation enables us to isolate and better understand the complementary roles and the relative importance of each component. Our experiments demonstrate that the combination of consistency- and confidence-based metrics outperforms the pure consistency-based measure on a vast majority of tasks. Notably, in the few cases where the pure consistency messure outperforms the combined approach, the performance difference is minimal. It is possible that a more suitable choice of a similarity measure or confidence-based metric for the task could further improve the performance.
This leads us to the next part of our ablation study, where we investigate the impact of different similarity measures (see Appendix B.2). We find that for some tasks, the similarity score computed by the Cross-encoder does not yield optimal performance. For example, for question-answering tasks on CoQA and Trivia, NLI-derived similarity performs better than the Cross-encoder similarity and outperforms the pure consistency-based uncertainty discussed above.
The next section of our ablation study focuses on alternative forms of combining model confidence $u_{*}^{\text{info}}$ and consistency $u_{*}^{\text{cons}}$ (see Apendix B.3). First, we consider an additive form of combining them: $\mathrm{U}_{\text{AdditiveCoCoA}}=u_{*}^{\text{info}}+u_{*}^{\text{cons}}$. The results show that this additive formulation does not perform as well compared to the multiplicative one. The additive form tends to underemphasize the interaction between the two components, which is critical for capturing the nuanced relationships between confidence and consistency.
We also consider an alternative formulation of the consistency term $u_{*}^{\text{cons}}$, as the average of the full pairwise dissimilarity. In this formulation, $u_{*}^{\text{cons}}$ represents the average inconsistency across all samples rather than focusing solely on the dissimilarity of the evaluated sequence with the other samples. Our experiments demonstrate that this formulation is not very strong. By distributing the consistency computation across all samples, it loses focus on the specific sequence being evaluated.
Lastly, in Appendix B.3, we also consider alternative formulations of the information-based metric that do not rely on logarithmic transformations. While we primarily use logarithms due to their numerical stability, we explore an alternative approach by converting these values back to probabilities and analyzing their impact on uncertainty estimation. Our findings indicate that both formulations exhibit consistent performance and yield similar results. This suggests that while logarithmic transformations enhance numerical stability, the choice between log-based and probability-based formulations does not affect much the overall performance.
5 Limitations
Task and Domain Dependency.
Our method relies on both an information-based confidence score and a semantic similarity function. The effectiveness of each can vary across models, tasks, and domains. For open-ended tasks with multiple equally valid outputs (e.g., creative generation), consistent rephrasing may inflate the perceived certainty. Conversely, in domains that demand highly precise factual or logical correctness (e.g., math problem solving), small deviations in reasoning can lead to large outcome differences that are not fully captured by a generic similarity measure. Adapting both the confidence measure and the similarity function to specific domains or prompt types is an important direction for future work.
Limited Sample Size.
CoCoA estimates the model’s consistency by sampling multiple outputs and comparing them. In practice, generating a large number of samples can be computationally expensive and may increase the inference latency. Consequently, our experiments (like many sampling-based approaches) rely on relatively small sample sets. Although even a handful of samples can provide a meaningful estimate of consistency, it may not fully capture the diversity of the underlying distribution for certain tasks or for more complex prompts.
Quality Metric.
Finally, the CoCoA’s performance assessment depends on quality metrics (e.g., COMET for machine translation, and Accuracy for QA) that may not capture every nuance of textual outputs. Automatic metrics can have blind spots, particularly in evaluating coherence, factual correctness, or subtle aspects of style. Further refining or extending quality metrics to account for deeper reasoning, factual faithfulness, and stylistic appropriateness would better align uncertainty scores with real-world perceptions of model correctness.
6 Conclusion
We presented CoCoA, a unified approach that integrates Confidence and Consistency for uncertainty quantification in LLMs. By combining token-level confidence scores with semantic similarity between multiple sampled outputs, CoCoA offers a more holistic view of uncertainty than either approach alone. In extensive evaluations on question answering, summarization, and translation, our approach outperformed existing baselines and state-of-the-art UQ methods. Moreover, CoCoA’s flexible design allows easy adaptation to a variety of tasks and settings.
Impact Statement
As large language models (LLMs) are increasingly integrated into everyday applications, ensuring the robustness of their outputs is essential, particularly in high-stakes domains such as healthcare and the legal sector. In addition to flagging low-quality responses, clear confidence estimates can help reduce overreliance on automated systems and foster safer, more responsible LLM usage. Here, we introduced CoCoA, a new framework that enhances uncertainty quantification in LLMs. While CoCoA improves reliability, its use requires careful domain adaptation to align with specific task requirements.
Appendix A Decoding Strategy and Sample Selection
To facilitate the choice of decoding and sample selection strategies for our experiments, we conducted an evaluation of model performance with different approaches to both. Table 3 shows average values of corresponding quality metrics for all combinations of models and datasets. We considered 4 approaches for the selection of output that subsequently is used to calculate the quality of generation.
  • Greedy decoding produces single output by selecting top-1 candidate token at each generation step, thus not further selection of sample is needed.
  • Best (normalized) sample selects the output with highest model-assigned (length-normalized) probability among several sampled outputs.
We note that selecting a random sample from the model outputs incurs a significant drop in the quality of results on several datasets, most prominently on GSM8k. Based on these observations, we evaluate the efficacy of UE on two setups: greedy decoding and stochastic sampling with a focus on the highest-probability sample.
In all experiments, we performed stochastic sampling with temperature $t=1.0$, top-k equal to 50, and top-p equal to $1.0$.
Dataset Metric Llama Mistral Falcon
Greedy Random Best Best Norm Greedy Random Best Best Norm Greedy Random Best Best Norm
Trivia Algin Score 0.686 0.625 0.687 0.689 0.743 0.655 0.750 0.751 0.557 0.473 0.568 0.562
MMLU Accuracy 0.590 0.430 0.597 0.587 0.633 0.558 0.632 0.632 0.713 0.639 0.713 0.713
CoQa Algin Score 0.499 0.359 0.529 0.462 0.574 0.403 0.591 0.528 0.512 0.408 0.560 0.471
GSM8k Accuracy 0.548 0.234 0.261 0.432 0.382 0.169 0.190 0.290 0.776 0.313 0.205 0.593
Xsum Align Score 0.848 0.608 0.825 0.830 0.803 0.578 0.775 0.777 0.842 0.734 0.782 0.809
WMT14FrEn Comet 0.863 0.819 0.852 0.859 0.863 0.812 0.830 0.855 0.867 0.833 0.857 0.863
WMT19DeEn Comet 0.870 0.816 0.854 0.860 0.864 0.805 0.836 0.851 0.846 0.807 0.826 0.838
Appendix B Ablation
B.1 Average Dissimilarity as Uncertainty Measure
Tables 4 and 5 report PRRs of CoCoA-family methods with uncertainty estimates based solely on average dissimilarity of samples, as proposed in equation (8) We observe that it is still widely beneficial to synthesize consistency of outputs with model confidence, even when limiting consistency evaluation to the particular sample to be scored.
@@ B.1 Average Dissimilarity as Uncertainty Measure
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
AveDissimilarity 0.051 0.285 0.500 0.379 0.647 0.423 0.435
$\text{CoCoA}_{MSP}$ 0.330 0.396 0.598 0.383 0.670 0.466 0.517
$\text{CoCoA}_{PPL}$ 0.286 0.375 0.568 0.369 0.674 0.466 0.467
$\text{CoCoA}_{MTE}$ 0.288 0.374 0.564 0.355 0.673 0.447 0.491
Llama8b-Base
AveDissimilarity 0.024 0.389 0.453 0.375 0.614 0.392 0.368
$\text{CoCoA}_{MSP}$ 0.378 0.456 0.582 0.349 0.597 0.485 0.372
$\text{CoCoA}_{PPL}$ 0.387 0.448 0.514 0.338 0.593 0.452 0.433
$\text{CoCoA}_{MTE}$ 0.380 0.446 0.511 0.337 0.601 0.402 0.447
Falcon7b-Base
AveDissimilarity 0.226 0.337 0.496 0.408 0.656 0.485 0.426
$\text{CoCoA}_{MSP}$ 0.257 0.433 0.578 0.396 0.684 0.529 0.436
$\text{CoCoA}_{PPL}$ 0.229 0.436 0.580 0.406 0.677 0.529 0.478
$\text{CoCoA}_{MTE}$ 0.228 0.439 0.577 0.395 0.685 0.517 0.510
Table 4: Comparison of PRRs of CoCoA-family methods with similarity of greedy output and other samples taken in isolation.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
Mistral7b-Base
AveDissimilarity 0.071 0.670 0.708 0.405 0.614 0.423 0.846
$\text{CoCoA}_{MSP}$ 0.366 0.712 0.730 0.430 0.644 0.466 0.562
$\text{CoCoA}_{PPL}$ 0.372 0.735 0.755 0.402 0.648 0.466 0.937
$\text{CoCoA}_{MTE}$ 0.373 0.708 0.732 0.373 0.645 0.447 0.935
Llama8b-Base
AveDissimilarity 0.030 0.473 0.598 0.395 0.600 0.353 0.795
$\text{CoCoA}_{MSP}$ 0.484 0.529 0.685 0.384 0.587 0.452 0.513
$\text{CoCoA}_{PPL}$ 0.450 0.544 0.689 0.364 0.573 0.422 0.925
$\text{CoCoA}_{MTE}$ 0.451 0.520 0.638 0.346 0.582 0.363 0.900
Falcon7b-Base
AveDissimilarity 0.282 0.491 0.651 0.416 0.627 0.484 0.979
$\text{CoCoA}_{MSP}$ 0.210 0.564 0.698 0.428 0.659 0.530 0.498
$\text{CoCoA}_{PPL}$ 0.310 0.579 0.717 0.415 0.644 0.530 1.000
$\text{CoCoA}_{MTE}$ 0.289 0.551 0.678 0.402 0.646 0.517 0.998
Table 5: Comparison of PRRs of CoCoA-family methods with similarity of samples with best sample taken in isolation.
B.2 Choice of Similarity Function
For sample consistency estimation, one could come up with a variety of similarity functions $g(\mathbf{y},\mathbf{y}^{\prime})$. We perform a comparison of the effectiveness of CoCoA-family methods using several such functions. We consider the following functions:
  • AlignScore (Zha et al., 2023) with AlignScore-large model
  • RougeL (Lin, 2004)
  • NLI (He et al., 2021) based on microsoft/deberta-large-mnli model
  • CrossEncoder (Liu et al., 2019) based on cross-encoder/stsb-roberta-large model.
Tables 6 and 7 report these results. There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand. We opt to report all results in other sections with CrossEncoder-based similarity as it by itself provides a significant improvement over baselines, and for consistency and ease of comparison reasons. However, when applying these methods to a particular task, we encourage users to select appropriate underlying similarity function for best results.
Method Dataset
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
@@ B.2 Choice of Similarity Function
Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with greedy sample taken in isolation.
@@
XSum WMT14FrEn WMT19DeEn CoQa Trivia MMLU GSM8k
@@
Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function with best sample taken in isolation.
B.3 Different Ways of Combining Confidence and Consistency
We justify the particular form of equation (9) by considering alternative ways to combine sample-focused confidence with consistency estimation. Results are presented in Tables 8 and 9. In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
$\mathrm{U}_{AdditiveCoCoA}=u_{*}^{\text{info}}+u_{*}^{\text{cons}},$ (11)

### v3

@@ Appendix B Results Summary for Most Probable Sample and MBR Sample strategies
MTE 0.476 0.469 0.449 0.559 0.637 0.350 0.602 0.492 0.186
$\text{CoCoA}_{MTE}$ 0.547 $\uparrow$ 0.579 $\uparrow$ 0.451 $\uparrow$ 0.600 $\uparrow$ 0.720 $\uparrow$ 0.373 $\uparrow$ 0.641 $\uparrow$ 0.614 $\uparrow$ 0.289 $\uparrow$
Table 3: Results for Evaluated Sequence – Most Probable Sample: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Metric Llama8b-Base Mistral7b-Base Falcon7b-Base
@@
MCSE 0.402 0.297 0.158 0.474 0.330 0.237 0.437 0.346 0.076
MCNSE 0.392 0.389 0.124 0.441 0.435 0.192 0.446 0.440 0.078
Semantic Entropy 0.443 0.316 0.159 0.507 0.352 0.236 0.458 0.373 0.077
DegMat 0.478 0.290 0.047 0.471 0.322 0.124 0.499 0.371 -0.018
EigValLaplacian 0.434 0.224 0.037 0.437 0.279 0.115 0.460 0.326 -0.022
SAR 0.509 0.456 0.187 0.532 0.448 0.266 0.531 0.478 0.083
Consistency 0.504 0.365 0.235 0.518 0.367 0.284 0.538 0.410 0.099
MSP 0.415 0.404 0.266 0.447 0.433 0.312 0.405 0.376 0.170
$\text{CoCoA}_{MSP}$ 0.527 $\uparrow$ 0.496 $\uparrow$ 0.293 $\uparrow$ 0.553 $\uparrow$ 0.515 $\uparrow$ 0.348 $\uparrow$ 0.500 $\uparrow$ 0.526 $\uparrow$ 0.154
PPL 0.408 0.421 0.164 0.432 0.452 0.273 0.445 0.438 0.100
$\text{CoCoA}_{PPL}$ 0.532 $\uparrow$ 0.504 $\uparrow$ 0.215 $\uparrow$ 0.554 $\uparrow$ 0.514 $\uparrow$ 0.328 $\uparrow$ 0.552 $\uparrow$ 0.537 $\uparrow$ 0.109 $\uparrow$
MTE 0.467 0.461 0.148 0.515 0.496 0.219 0.490 0.521 0.069
$\text{CoCoA}_{MTE}$ 0.542 $\uparrow$ 0.479 $\uparrow$ 0.201 $\uparrow$ 0.574 $\uparrow$ 0.486 0.290 $\uparrow$ 0.562 $\uparrow$ 0.527 $\uparrow$ 0.081 $\uparrow$
Table 4: Results for Evaluated Sequence - MBR Decoding: Mean PRR across datasets for each task. The best performing method is in bold, and the second-best is underscored. Arrows indicate improvement in CoCoA over the base version.
Appendix C Ablation
C.1 Choice of Similarity Function
For sample consistency estimation, one could come up with a variety of similarity functions $s(\mathbf{y},\mathbf{y}^{\prime})$. We perform a comparison of the effectiveness of CoCoA-family methods using several such functions. We consider the following functions:
  • AlignScore Zha et al. [2023] with AlignScore-large model;
  • RougeL Lin [2004];
  • NLI He et al. [2021] based on microsoft/deberta-large-mnli model;
  • CrossEncoder Liu et al. [2019] based on cross-encoder/stsb-roberta-large model.
Tables 5, 6 and 7 report these results. There exists a considerable variation of relative effectiveness of proposed methods with various similarity function choices, depending on a task at hand. We opt to report all results in other sections with CrossEncoder-based similarity as it by itself provides a significant improvement over baselines, and for consistency and ease of comparison reasons. However, when applying these methods to a particular task, we encourage users to select appropriate underlying similarity function for best results.
@@ C.1 Choice of Similarity Function
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
@@
Table 5: Comparison of PRRs of CoCoA-family methods with different choices of the similarity function. Main model response obtained by greedy decoding.
@@
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
@@
Table 6: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by selecting the most probable sample.
Method Dataset
XSUM WMT14FrEn WMT19DeEn CoQA TriviaQA MMLU GSM8k
Mistral7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.287 0.326 0.460 0.312 0.575 0.515 0.747
RougeL 0.259 0.420 0.562 0.304 0.502 0.515 0.750
NLI 0.278 0.375 0.506 0.350 0.589 0.512 0.766
CrossEncoder 0.295 0.441 0.589 0.336 0.597 0.514 0.766
$\text{CoCoA}_{PPL}$
AlignScore 0.283 0.333 0.482 0.330 0.596 0.515 0.738
RougeL 0.219 0.427 0.561 0.300 0.517 0.515 0.744
NLI 0.219 0.390 0.508 0.334 0.602 0.512 0.723
CrossEncoder 0.258 0.451 0.577 0.322 0.612 0.514 0.768
$\text{CoCoA}_{MTE}$
AlignScore 0.269 0.314 0.447 0.357 0.663 0.497 0.746
RougeL 0.181 0.446 0.571 0.327 0.622 0.497 0.786
NLI 0.182 0.370 0.482 0.348 0.686 0.499 0.704
CrossEncoder 0.236 0.419 0.553 0.337 0.680 0.494 0.783
Llama8b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.221 0.309 0.462 0.261 0.467 0.622 0.717
RougeL 0.195 0.401 0.542 0.250 0.361 0.631 0.721
NLI 0.222 0.311 0.463 0.311 0.480 0.580 0.769
CrossEncoder 0.225 0.403 0.590 0.304 0.487 0.586 0.731
$\text{CoCoA}_{PPL}$
AlignScore 0.225 0.325 0.484 0.304 0.476 0.603 0.747
RougeL 0.169 0.416 0.537 0.274 0.370 0.612 0.758
NLI 0.180 0.334 0.460 0.319 0.488 0.558 0.761
CrossEncoder 0.198 0.428 0.580 0.320 0.496 0.562 0.751
$\text{CoCoA}_{MTE}$
AlignScore 0.209 0.293 0.460 0.320 0.563 0.556 0.749
RougeL 0.148 0.436 0.559 0.314 0.497 0.587 0.785
NLI 0.156 0.298 0.445 0.340 0.592 0.460 0.740
CrossEncoder 0.189 0.398 0.561 0.345 0.587 0.472 0.765
Falcon7b-Base
$\text{CoCoA}_{MSP}$
AlignScore 0.279 0.307 0.473 0.312 0.594 0.574 0.460
RougeL 0.202 0.424 0.487 0.307 0.566 0.574 0.447
NLI 0.231 0.362 0.508 0.355 0.618 0.577 0.531
CrossEncoder 0.248 0.461 0.590 0.336 0.617 0.575 0.472
$\text{CoCoA}_{PPL}$
AlignScore 0.278 0.320 0.534 0.350 0.607 0.574 0.617
RougeL 0.190 0.428 0.556 0.333 0.571 0.574 0.632
NLI 0.216 0.376 0.549 0.368 0.623 0.578 0.661
CrossEncoder 0.242 0.462 0.613 0.345 0.624 0.575 0.662
$\text{CoCoA}_{MTE}$
AlignScore 0.277 0.302 0.522 0.382 0.657 0.556 0.598
RougeL 0.175 0.472 0.573 0.364 0.642 0.556 0.617
NLI 0.196 0.359 0.530 0.400 0.671 0.562 0.639
CrossEncoder 0.237 0.454 0.600 0.373 0.666 0.557 0.650
Table 7: Comparison of PRRs of CoCoA-family methods with different choices of similarity function. Main model response obtained by MBR decoding.
C.2 Different Ways of Combining Confidence and Consistency
We justify the particular form of equation (11) by considering alternative ways to combine sample-focused confidence with consistency estimation. Results are presented in Tables 89 and 10. In particular, we investigate the performance of the additive approach (AdditiveCoCoA):
$U_{\text{AdditiveCoCoA}}(\mathbf{y}_{*}\mid\mathbf{x})=u(\mathbf{y}_{*}\mid\mathbf{x})+\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x}),$ (16)